On moving a charge of 20 coulomb by 2 cm, 2 J of work is done, then the potential difference between the points is
(a) 0.1 V
(b) 8V
(c) 2V
(d) 0.5V

Grade:6

2 Answers

Navjyot Kalra
9 years ago

(a) We know that WAB/q = VB - VA

? VB – VA = 2 J /20 C = 0.1 J/C = 0.1 V

5. (a) For an isolated sphere, the capacitance is given by

C = 4p ?0 r = 1/ 9x 109 x 1 = 1.1 x 10-10 F
vanshika saini
5 years ago
the total flux across the sphere of a guassian sphere encloses an electric dipole will be............